Solve for the missing information designated by “?” in the following table. (Use 365 days in a year. Round the inventory turnover ratio to one decimal place before computing days to sell. Round days to sell to one decimal place.)BI CGS E L Inventory Turnover Ratio Days to Sell Case $ | $ Purchases $ 1,160 260 250 $ $ $ 1,080 1,840 1,640 9.2 $ 70 22.3
Answer #1
Case BI Purchases CGS EI Inventory turnover ratio Days to sell
a $260 $1160 $1080 $340 3.6 101.4
b $250 $1740 $1840 $150 9.2 39.7
c. $130 $1580 $1640 $70 16.4 22.3

Case a

Ending inventory= Beginning inventory+Purchases-Cost of goods sold

= $260+1160-1080= $340

Average inventory= (Beginning inventory+Ending inventory)/2= ($260+340)/2= $300

Inventory turnover ratio= Cost of goods sold/Average inventory

= $1080/300= 3.6

Days to sell= 365 days/Inventory turnover ratio= 365/3.6= 101.4 days

Case b

Inventory turnover ratio= Cost of goods sold/Average inventory

9.2= $1840/Average inventory

Average inventory= $1840/9.2= $200

Average inventory= (Beginning inventory+Ending inventory)/2

$200= (250+Ending inventory)/2

$400 = 250+Ending inventory

Ending inventory= $150

Purchases= Cost of goods sold-Beginning inventory+Ending inventory

= $1840-250+150= $1740

Days to sell= 365 days/Inventory turnover ratio= 365/9.2= 39.7

Case c

Days to sell= 365 days/Inventory turnover ratio

22.3= 365/Inventory turnover ratio

Inventory turnover ratio= 365/22.3= 16.4

Inventory turnover ratio= Cost of goods sold/Average inventory

16.4= $1640/Average inventory

Average inventory= $1640/16.4= $100

Average inventory= (Beginning inventory+Ending inventory)/2

$100= (Beginning inventory+70)/2

$200= Beginning inventory+70

Beginning inventory= $130

Purchases= Cost of goods sold-Beginning inventory+e=Ending inventory

= $1640-130+70= $158w0